OpenStudy (anonymous):

what are the solutions of 3x^2 -22x = -24?

OpenStudy (anonymous):

(3x-4)(x-6) = 0 \[x=\frac{ 4 }{ 3}\]\[x=6\]\[3x(x)-18x-4x+24=0\]\[ 3x^2-22x-24=0\]\[3x^2-22x=-24\]

OpenStudy (anonymous):

thank you so much